# M3-20M: A Large-Scale Multi-Modal Molecule Dataset for AI-driven Drug Design and Discovery

### v1

@@ 3.3 Generated Text Quality Control by Human Expert Scoring
To ensure the reliability and utility of the generated textual descriptions, we implement an expert scoring mechanism inspired by some recent works 50, 51. We recruit six undergraduate and postgraduate chemistry students from top China universities, all having excellent academic records. They evaluate the accuracy and usefulness of the generated texts, ensuring that the textual descriptions are both scientifically valid and practically useful.
This mechanism assesses the textual descriptions from four dimensions: accuracy, effectiveness, comprehensiveness, and simplicity, as established in previous works on natural language generation and evaluation 52, 53. Each description can earn a maximum score of 10 points. The scoring system is structured as follows:
@@
4 Statistic Results
This section presents some statistic information of the M3-20M dataset to highlight its comprehensive annotations, diverse textual descriptions and physicochemical properties, which can significantly enhance its utility for molecular research and AI-driven drug discovery. By comparing it with existing datasets such as PubChem, we emphasize the enhancements in both scale and quality that M3-20M offers. The section further delves into word distribution — a key feature of textual data, and explores correlations among physicochemical properties through advanced statistical techniques like Principal Component Analysis (PCA). Additionally, we outline the robust maintenance protocols established to ensure our dataset’s accuracy and usability for the research community. These findings demonstrate the dataset’s value as a fundamental resource for molecular representation learning and related applications.
4.1 Textual Descriptions
Figure 4 shows a comparison of molecule descriptions between PubChem and our M3-20M. The PubChem description bar is relatively small, reaching only 360,133, while the M3-20M annotation description bar towers over it, marking a count of 20,249,090. This enhancement is critical for AI-driven drug design and discovery, as the increased volume and details of descriptions facilitate more comprehensive analysis and model training.
Figure 4: Comparison of molecule descriptions: PubChem vs. our M3-20M
4.2 Word Distribution in Description Texts
Figure 5 presents the top 20 most frequently occurring words in the textual descriptions of our dataset, plotted alongside their respective frequencies to emphasize their relative prominence. The most commonly used terms, such as ’natural,’ ’product,’ and ’acid,’ appear tens of thousands of times, underscoring their central role in describing molecular properties and functions. Additional notable terms, including ’related,’ ’organisms,’ and ’amino,’ further highlight the biological and chemical contexts associated with these molecules. This visualization of word frequencies provides valuable insights into molecular datasets’ linguistic patterns and descriptive focus.
Figure 5: Top 20 most frequent words in textual descriptions.
4.3 Physicochemical Properties
Table 2 and Figure 6 provide a detailed overview of the specific names, corresponding quantities, and reference of the 26 key physicochemical properties extracted from PubChem. Of the 26 key physicochemical properties, 18 are computed using professional software  [^1] PubChem 2.2; Cactvs 3.4.8.18, while the remaining 8 represent experimental values sourced from the CAMEO Chemicals database and Hazardous Substances Data Bank (HSDB). Notably, 18 of these properties exhibit quantities reaching tens of millions, underscoring our dataset’s extensive coverage and comprehensiveness.
Physicochemical Properties Number Reference Molecular Weight 19174000 Computed by PubChem 2.2 XLogP3 16340203 Computed by XLogP3 3.0 Hydrogen Bond Donor Count 19173687 Computed by Cactvs 3.4.8.18 Hydrogen Bond Acceptor Count 19173559 Computed by Cactvs 3.4.8.18 Rotatable Bond Count 19173464 Computed by Cactvs 3.4.8.18 Exact Mass 19173398 Computed by PubChem 2.2 Monoisotopic Mass 19173327 Computed by PubChem 2.2 Topological Polar Surface Area 19173287 Computed by Cactvs 3.4.8.18 Heavy Atom Count 19173259 Computed by PubChem Formal Charge 19173222 Computed by PubChem Complexity 19173210 Computed by Cactvs 3.4.8.18 Isotope Atom Count 19173193 Computed by PubChem Defined Atom Stereocenter Count 19173170 Computed by PubChem Undefined Atom Stereocenter Count 19173153 Computed by PubChem Defined Bond Stereocenter Count 19173147 Computed by PubChem Undefined Bond Stereocenter Count 19173137 Computed by PubChem Covalently-Bonded Unit Count 19173132 Computed by PubChem Compound Is Canonicalized 19173130 Computed by PubChem Physical Description 2031 CAMEO Chemicals database Color/Form 246 Hazardous Substances Data Bank (HSDB) Odor 100 Hazardous Substances Data Bank (HSDB) Boiling Point 130 CAMEO Chemicals database Melting Point 633 CAMEO Chemicals database Flash Point 35 CAMEO Chemicals database Solubility 10813 CAMEO Chemicals database Density 170 CAMEO Chemicals database
Figure 6: Physicochemical property distribution.
Figure 7: Correlation matrix of physicochemical properties.
Figure 8: Visualization of molecular space with principal component analysis (PCA).
As shown in Figure 7, we first calculate the correlation matrix of all physicochemical properties and draw a heat map. We find that Molecular Weight and XLogP3 are the most suitable because they exhibit strong correlations with several other key properties, indicating their ability to capture significant variance in the dataset. Thus, we select these two properties as the first and second principal components for Principal Components Analysis (PCA) and visualize the molecular space of randomly sampled 10000 molecules as shown in Figure 8.
4.4 Dataset Maintenance
To ensure the accessibility and accuracy of the M3-20M dataset, we have organized a dedicated maintenance team consisting of six members, including three doctoral students and three master’s students. This team will perform regular maintenance twice a week, correcting any errors and updating the dataset as needed. This rigorous maintenance protocol supports the community’s exploration of molecular representation and downstream molecular tasks and facilitates the training of large models with improved versatility and robustness.
4.5 Supporting Various Downstream Tasks
The M3-20M dataset supports model training and tuning for various downstream drug design and discovery tasks. While general large language models have achieved remarkable success in natural language processing, their performance in specialized fields like drug design and discovery still has significant room for improvement. Figure 9 illustrates how to use the M3-20M dataset to enhance large language models (LLMs) or train models for downstream drug design and discovery tasks, which consists of three paradigms: prompting, finetuning, and training from scratch.
Prompt engineering involves carefully constructing input prompts to guide a general large language model (LLM) in generating more accurate outputs tailored to drug design and discovery. By leveraging our M3-20M dataset, prompt engineering allows the model to quickly adapt to this field without altering its internal parameters. Fine-tuning refers to further training the LLMs on the M3-20M dataset, adjusting certain model parameters based on existing pre-training. This process enables the model to better align with the specific characteristics and needs of drug design and discovery while retaining its general language knowledge. Training from scratch entails building and training a language model on drug design and discovery domain data. This approach necessitates a large volume of training data, which our M3-20M dataset can adequately provide. By utilizing our M3-20M dataset through prompt engineering, fine-tuning, and training from scratch, the general LLM’s drug design and discovery capabilities can be enhanced from multiple perspectives and to varying degrees.
M3-20M also can aid in tasks such as pharmacokinetics modeling and drug-target interaction prediction. Moreover, we construct seven multimodal sub-datasets, QM9-MM, MOSES-MM, BACE-MM, BBBP-MM, HIV-MM, ClinTox-MM, and Tox21-MM, to facilitate more accurate and robust molecular property prediction. We also provide tools for generating molecular images to support compound patent searches and for crawling PubMed to find articles related to specific molecules, thereby enhancing textual descriptions.
@@ 4.5 Supporting Various Downstream Tasks
5 Dataset Evaluation
In this section, we evaluate the effectiveness of the M3-20M dataset on two fundamental downstream tasks: molecule generation and molecular property prediction. We primarily perform prompt tuning and in-context learning on the closed-source large language models GLM4, GPT-3.5, and GPT-4, while also conducting fine-tuning on the open-source model Llama3-8b. Additionally, our dataset can be applied to other tasks, including name prediction, reaction prediction, retrosynthesis, text-based molecule design, and molecule captioning. Our findings demonstrate that M3-20M significantly enhances model performance across these molecular tasks.
5.1 Molecule Generation
We utilize the GLM-4, GPT-3.5, and GPT-4 APIs to evaluate the power of M3-20M in molecule generation. Through in-context learning (ICL) 54, we provide 10 randomly selected multi-modal examples (including SMILES strings of molecules, three-dimensional atomic coordinates, and descriptions) and single-modal examples (containing only SMILES), respectively. Subsequently, 100 molecules are generated randomly for each setup.
We analyze the performance of models trained on single-modal and multi-modal datasets in molecule generation using three metrics: validity, uniqueness, and novelty. The results, presented in Table 3, demonstrate the superiority of multi-modal datasets across all metrics and models.
Validity measures the proportion of chemically valid generated molecules. Multi-modal datasets consistently outperform single-modal ones. For instance, GLM4’s validity improves from 72.73% to 85.37%, GPT-3.5’s from 76.01% to 84.8%, and GPT-4’s from 92.3% to 97.99%.
Uniqueness assesses the diversity of generated molecules. Multi-modal datasets show significant improvements: GLM4’s uniqueness rises from 68.17% to 82.25%, GPT-3.5’s from 46.46% to 93.86%, and GPT-4’s from 58.64% to 70.17%.
Novelty evaluates the proportion of generated novel molecules. Multi-modal datasets show substantial improvements, with GLM4’s novelty increasing from 97.90% to 98.10%, GPT-3.5’s from 85.72% to 96.51%, and GPT-4’s from 90.66% to 98.95%.
@@ 5.1 Molecule Generation
Furthermore, Table 4 summarizes the molecular generation performance of our M3-20M dataset using MOSES-MM as the benchmark. Multi-modal data enhances SNN/Test performance for GPT-3.5 and GPT-4, showing increased similarity to the test set. For Frag/Test, multi-modal data boosts GLM4 to 0.27% and GPT-4 to 0.20%. In Scaf/Test, all models improve with multi-modal data; GLM4 rises to 0.07%, while GPT-3.5 and GPT-4 each improve to 0.01%. Regarding Filters, GLM4’s pass rate increases to 0.80% with multi-modal data, while GPT-3.5 maintains 1.0%.
Metrics like logP, SA, QED, and molecular weight are assessed using the Wasserstein-1 distance to the MOSES-MM test set, where lower values indicate better performance. Multi-modal data significantly reduces logP distance, especially for GPT-3.5 (0.99) and GPT-4 (0.92), indicating superior predictions. In SA, multi-modal data improves scores for GPT-3.5 and GPT-4, reducing distances to 0.57 and 0.54, respectively. QED scores improve with multi-modal data, reaching 0.12 for both models. For molecular weight, predictions improve for GLM4, reducing the distance to 105.72.
These findings demonstrate that our multi-modal dataset significantly enhances molecule generation models across nearly all metrics, highlighting the benefit of integrating diverse molecular representations. The results support the widespread use of multi-modal data in molecular research to develop more robust and versatile generative models.
GLM4 GPT-3.5 GPT-4 SNN/Test(%)↑ Single-modal 0.49(±0.03) 0.53(±0.09) 0.49(±0.07) Multi-modal 0.44(±0.09) 0.56(±0.04) 0.52(±0.01) Frag/Test(%)↑ Single-modal 0.24(±0.02) 0.44(±0.12) 0.19(±0.08) Multi-modal 0.27(±0.14) 0.32(±0.01) 0.20(±0.01) Scaf/Test(%)↑ Single-modal 0.04(±0.05) 6.90e-4(±6.90e-4) 2.01e-3(±4.13e-4) Multi-modal 0.07(±0.09) 0.01(±0.01) 0.01(±4.99e-3) Filters(%)↑ Single-modal 0.73(±0.09) 0.8(±0.0) 1.0(±0.0) Multi-modal 0.80(±1.11e-16) 1.0(±0.0) 0.9(±0.10) logP(↓) Single-modal 1.35(±0.17) 1.63(±0.32) 1.27(±0.31) Multi-modal 1.44(±0.28) 0.99(±0.05) 0.92(±0.11) SA(↓) Single-modal 0.83(±0.22) 0.63(±0.11) 0.60(±0.14) Multi-modal 1.00(±0.10) 0.57(±0.02) 0.54(±0.06) QED(↓) Single-modal 0.20(±0.04) 0.15(±0.05) 0.15(±0.06) Multi-modal 0.20(±0.08) 0.12(±0.03) 0.12(±0.02) weight(↓) Single-modal 112.10(±20.66) 67.83(±15.92) 89.75(±37.98) Multi-modal 105.72(±38.35) 68.01(±4.04) 95.62(±9.33)
Table 4: Molecule generation performance comparison on MOSES-MM.
@@
QM9 48 75.17 (±1.43) 64.13 (±1.58) 89.43 (±0.01)
Pcdes 29 84.31 (±0.24) 82.64 (±1.60) 93.253 (±0.59)
PubChemSTM 28 82.24 (±0.70) 80.89 (±1.12) 95.91 (±1.93)
igcdata 23 78.34 (±2.22) 61.21 (±1.50) 95.07 (±0.07)
@@
Table 5: Model performance comparison among different datasets.
Finally, we compare the molecule generation performance of GPT-3.5 across various datasets: QM9, Pcdes, PubChemSTM, igcdata, and our M3-20M, focusing on Validity, Uniqueness, and Novelty. Table 5 presents the experimental results.
Using our dataset, the model achieves the highest validity score of 84.8%, surpassing all other datasets, indicating the effectiveness of M3-20M in generating chemically accurate molecules. For uniqueness, our dataset again leads with 93.86%, significantly higher than Pcdes and PubChemSTM, demonstrating its capacity to support the generation of diverse molecular structures essential for discovering novel compounds. Our dataset also achieves the highest novelty score of 96.51%, followed by PubChemSTM at 95.91% and Pcdes at 93.25%, underscoring its effectiveness in generating unique molecular structures.
Overall, the comparison across Validity, Uniqueness, and Novelty highlights the superiority of our M3-20M dataset. It enables the generation of chemically valid molecules with high diversity and novelty, making it a valuable resource for model training and tuning in drug design and discovery.
5.2 Molecular Property Prediction
We refer to MoleculeNet 4 and divide the molecular property prediction task into regression task and classification task.
Regression tasks:
To test the effectiveness of our multi-modal dataset in support of molecule property prediction tasks, we conduct an ablation experiment on the QM9-MM dataset with three settings: using only the SMILES strings of molecules; using both the SMILES strings and 3D coordinates of molecules; and using the SMILES strings, 3D coordinates and the textual descriptions of molecules. We consider two regression tasks: predicting the dipole moment ($\mu$) and isotropic polarizability ($\alpha$) of the given molecules. We employ GPT-3.5, GPT-4, and GLM-4 APIs to infer the molecular properties respectively. For each molecule to be inferred, we provide the corresponding molecule information in the prompt based on the different versions of the settings and randomly select four additional molecule examples to help the LLMs better understand the chemical structure and the relation between the molecule information and the prediction ground truth.
The metric we use for the regression tasks is Mean Absolute Error (MAE), measuring the average magnitude of errors between predicted values and actual values. The smaller the MAE value, the more accurate the prediction. The formula of the MAE metric is as follows:

### v2

@@ 3.3 Generated Text Quality Control by Human Expert Scoring
To ensure high-quality textual descriptions, we design a quality control protocol inspired by some recent works mccann2018natural ; kwiatkowski2019natural , which consists of three parts: selection and training of evaluation experts (or validators), scoring criteria, and stratified random sampling.
3.3.1 Evaluation Expert Selection and Training
To evaluate the quality of generated texts, six chemistry specialists (3 undergraduate and 3 postgraduate students) from top Chinese universities with documented expertise in molecular characterization were recruited as validators. All validators received a standardized training program covering evaluation criteria, error classification, and scoring calibration using 200 pilot samples.
3.3.2 Quality Assessment Criteria
This mechanism assesses the textual descriptions from four dimensions: accuracy, effectiveness, comprehensiveness, and simplicity, as established in previous works on natural language generation and evaluation wang2018glue ; post2018call . Each description can earn a maximum score of 10 points. The scoring system is structured as follows:
@@ 3.3.2 Quality Assessment Criteria
3.3.3 Sampling Strategy and Validation Results
We adopt stratified random sampling to sample 12,000 molecules (2,000 per validator) from our dataset, ensuring no overlap between validators. The validation reveals that the qualification rate is 98.8% (11,856/12,000), i.e., with only 144 instances requiring regeneration (24 ± 3.1 per validator). All the regenerated text descriptions underwent a secondary validation. This rigorous validation ensures that $\geq$ 99% of released descriptions meet strict scientific standards.
3.4 Multi-Modal Datasets for Downstream Tasks
To support model training for downstream tasks, we construct multi-modal datasets by integrating PubChem CIDs and descriptive text, thereby providing richer semantic context and molecular identity. In particular, we derive seven additional datasets from M3-20M, they are MOSES-MM and QM9-MM for molecule generation, while QM9-MM, BBBP-MM, BACE-MM, HIV-MM, Tox21-MM, and ClinTox-MM for molecular property prediction. Table 2 summarizes these datasets.
Table 2: Statistics of the multi-modal datasets this work augmented for downstream tasks, including the number of tasks, the number of molecules, task type, and evaluation metrics
Dataset #Tasks #Molecules Task type Metric
MOSES-MM 1 1,936,962 Generation SNN/Test, Frag/Test, Scaf/Test, Filters, logP, SA, QED, weight
QM9-MM 2 6,830 Generation and Regression Validity, Uniqueness, Novelty, MAE
BBBP-MM 1 2,039 Classification ROC–AUC
BACE-MM 1 1,513 Classification ROC–AUC
ClinTox-MM 2 1,478 Classification ROC–AUC
Tox21-MM 12 7,831 Classification ROC–AUC
HIV-MM 1 41,127 Classification ROC–AUC
4 Experiments
In this section, we evaluate the value and potential of the M3-20M dataset by experiments of two fundamental downstream tasks: molecule generation and molecular property prediction. In the experiments, we employ two different model adaptation approaches: (1) Few-shot prompting with in-context learning (FP-ICL) brown2020language on closed-source models (GLM4, GPT-3.5, and GPT-4), where we design input prompts with demonstration examples from M3-20M to guide the models’ responses without modifying model parameters; (2) Parameter-efficient fine-tuning using low-rank adaptation of large language models (LoRA) hu2021lora on the open-source Llama3-8b model, which updates a small subset of model parameters while maintaining the base model’s knowledge. Additionally, we present comprehensive statistical analyses on M3-20M to demonstrate its potential for supporting downstream drug design and discovery tasks. Our findings show that M3-20M can significantly improve model performance in the two molecular tasks with both adaptation approaches.
4.1 LLM Adaptation Methods
Large language models (LLMs) can be adapted for domain-specific tasks through various methods. As illustrated in Fig. 4, we employ two methods that differ significantly in their resource requirements and implementation.
4.1.1 Few-shot Prompting with In-context Learning (FP-ICL)
This approach involves providing carefully selected example pairs from M3-20M alongside the target task in a single prompt. By incorporating these demonstrations, the model learns to recognize patterns and generate appropriate responses for new cases without any parameter update. This method requires only API access to LLMs (such as GPT-4, GPT-3.5, and GLM4) and minimal computational resources, making it highly accessible and cost-effective for practical applications.
4.1.2 Parameter-efficient Fine-tuning with LoRA
Low-Rank Adaptation (LoRA) represents a more sophisticated adaptation method that modifies a subset of model parameters while preserving the model’s base knowledge. This approach requires access to open-source models (such as Llama3-8b) and substantial computational resources (e.g. RTX4090 GPU). While more resource-intensive, LoRA enables deeper model adaptation for specialized tasks.
While M3-20M is also suitable for training LLMs from scratch, such training requires powerful computational infrastructure beyond the scope of this study. Our adoption of these two adaptation methods provides practical insights into the dataset’s utility under different resource constraints.
4.2 Molecule Generation Experiments
The M3-20M dataset enables de novo molecule generation by providing rich contextual information that improves the understanding of chemical structures and properties of generative models.
4.2.1 Experimental Details
We employ the FP-ICL method to leverage multi-modal information for molecular generation by LLMs. Both general-purpose closed-source LLMs (GLM-4, GPT-3.5, GPT-4) and specialized generative models (MoFlow and D2L-OMP) are used. We compare the performance of using multi-modal molecules and single-modal molecules as examples for generation. For the multi-modal examples, we provide 10 randomly selected molecules from our M3-20M dataset with their SMILES strings, 3D coordinates, and textual descriptions to the LLMs. The single-modal examples use only SMILES strings. 100 molecules per configuration are generated for performance assessment. We trained these models using QM9-MM to maintain compatibility with their existing network structures and parameters. Both models are trained on RTX4090 with 24GB memory and batch size is 32. The flow-based model (MoFlow) is trained for 200 epochs, while the diffusion-based model (D2L-OMP) is trained for 300 epochs. The rest of the parameters of the two models are the same as those in the original paper. We perform each experiment five times to ensure statistical reliability and report the averaged results and variance.
We employ the FP-ICL method to leverage multi-modal information for molecular generation by LLMs. Both general-purpose closed-source LLMs (GLM-4, GPT-3.5, GPT-4) and specialized generative models (MoFlow and D2L-OMP) are used. We compare the performance of using multi-modal molecules and single-modal molecules as examples for generation. For the multi-modal examples, we provide 10 randomly selected molecules from our M3-20M dataset with their SMILES strings, 3D coordinates, and textual descriptions to the LLMs. The single-modal examples use only SMILES strings. 100 molecules per configuration are generated for performance assessment. We train MoFlow and D2L-OMP using QM9-MM on RTX4090 with 24GB memory with the batch size being 32. MoFlow is trained with 200 epochs, while D2L-OMP is trained with 300 epochs. The other parameters of these two models are the same as those in the original paper. We perform each experiment five times to ensure statistical reliability and report the averaged results and variance.
@@ 4.2.1 Experimental Details
4.2.2 Performance Metrics
We analyze the performance of models trained on single-modal and multi-modal datasets in molecule generation using three metrics: validity, uniqueness, and novelty.
Validity measures the proportion of chemically correct molecular structures generated. Uniqueness reflects the diversity of generated molecules. Novelty is defined as the fraction of valid molecules that are not present in the training dataset, revealing the model’s ability to propose new molecular structures.
Furthermore, we validate our results using the comprehensive MOSES benchmark with additional structural metrics. Similarity to the nearest neighbor (SNN/Test) measures reference manifold matching precision. Fragment similarity (Frag/Test) assesses BRICS fragment degen2008art distribution alignment in generated and reference sets. Scaffold similarity (Scaff/Test) evaluates Bemis-Murcko scaffold bemis1996properties distribution correspondence in generated and reference sets. Filters is the fraction of generated molecules that pass custom medicinal chemistry filters (MCFs) and PAINS filters baell2010new .
The MOSES benchmark includes key chemical property metrics (logP, SA, QED, Molecular weight (MW)). We calculate Wasserstein-1 distances between generated and test set distributions for these properties. Lower distances indicate better alignment with realistic molecular characteristics, further confirming generation quality.
4.2.3 Experimental Results
The results in Table 3 demonstrate that models leveraging multi-modal datasets consistently outperform their single-modal counterparts using single-modal datasets across all metrics and model types.
For validity, both specialized molecule generation models and general LLMs show remarkable performance improvement with multi-modal data.
For uniqueness, the consistently high uniqueness scores across all models demonstrate multi-modal datasets boost their robust capability to generate diverse molecular structures.
For novelty, our multi-modal dataset significantly enhances the performance of both LLMs and specialized molecule generation models.
@@ 4.2.3 Experimental Results
Specialized Molecule Generation Models
MoFlow 95.36($\pm$1.2) 96.21($\pm$1.24) 98.65($\pm$0.6) 98.71($\pm$0.19) 94.72($\pm$0.8) 99.12($\pm$0.13)
D2L-OMP 98.60 ($\pm$0.83) 98.88($\pm$0.92) 99.80($\pm$0.17) 99.80($\pm$0.14) 83.39($\pm$0.52) 89.10($\pm$0.51)
General LLMs
@@
Furthermore, Table 4 summarizes the results of additional molecule generation performance metrics. Multi-modal training consistently enhances performance across nearly all indicators, enabling the model to exceed single-modal baselines in generating chemically meaningful structures. Improvements are evident in SNN, Frag, Scaff, and Filter scores. Multi-modal training also reduces Wasserstein-1 distance for logP, SA, and QED properties. These results demonstrate that the chemical properties of generated molecules align with those of the reference molecules, highlighting the effectiveness of multi-modal data in producing compounds with pharmacologically relevant characteristics while maintaining structural diversity and synthetic accessibility.
Table 4: Molecule generation comparison using MOSES benchmark metrics. The table presents results of structural validity (SNN/Test, Frag/Test, Scaff/Test, Filters) and chemical property alignment (Wasserstein-1 distances for logP, SA, QED, MW) between using single-modal data and using multi-modal data. Lower Wasserstein-1 distance indicates better distribution matching with the reference compounds.
GLM4 GPT-3.5 GPT-4 SNN/Test(↑) Single-modal 0.49(±0.03) 0.53(±0.09) 0.49(±0.07) Multi-modal 0.49(±0.09) 0.56(±0.04) 0.52(±0.01) Frag/Test(↑) Single-modal 0.24(±0.02) 0.32(±0.12) 0.19(±0.08) Multi-modal 0.27(±0.14) 0.44(±0.01) 0.20(±0.01) Scaf/Test(↑) Single-modal 0.04(±0.05) 6.90e-4(±6.90e-4) 2.01e-3(±4.13e-4) Multi-modal 0.07(±0.09) 0.01(±0.01) 0.01(±4.99e-3) Filters(↑) Single-modal 0.73(±0.09) 0.8(±0.0) 1.0(±0.0) Multi-modal 0.80(±1.11e-16) 1.0(±0.0) 1.0(±0.10) logP(↓) Single-modal 1.35(±0.17) 1.63(±0.32) 1.27(±0.31) Multi-modal 1.44(±0.28) 0.99(±0.05) 0.92(±0.11) SA(↓) Single-modal 0.83(±0.22) 0.63(±0.11) 0.60(±0.14) Multi-modal 1.00(±0.10) 0.57(±0.02) 0.54(±0.06) QED(↓) Single-modal 0.20(±0.04) 0.15(±0.05) 0.15(±0.06) Multi-modal 0.20(±0.08) 0.12(±0.03) 0.12(±0.02) Molecular weight(↓) Single-modal 112.10(±20.66) 67.83(±15.92) 89.75(±37.98) Multi-modal 105.72(±38.35) 68.01(±4.04) 95.62(±9.33)
Table 5: Model performance comparison among different datasets.
@@
QM9 ramakrishnan2014quantum 75.17 (±1.43) 64.13 (±1.58) 89.43 (±0.01)
Pcdes zeng2022deep 84.31 (±0.24) 82.64 (±1.60) 93.253 (±0.59)
PubChemSTM liu2023multi 82.24 (±0.70) 80.89 (±1.12) 95.91 (±1.93)
igcdata liu2024git 78.34 (±2.22) 61.21 (±1.50) 95.07 (±0.07)
@@
Finally, we compare the molecule generation performance of GPT-3.5 using various datasets—QM9, Pcdes, PubChemSTM, igcdata, and our M3-20M—focusing on three key metrics: Validity, Uniqueness, and Novelty (see Table 5). The results demonstrate the advantages of our M3-20M dataset.
For validity, GPT-3.5 consistently generates the most chemically accurate molecules using our dataset, underscoring M3-20M’s effectiveness in capturing essential chemical properties. Regarding uniqueness, the molecules produced with our data exhibit a greater diversity of structures, indicating the potential for discovering novel compounds. This trend continues with the novelty metric, that is, using our dataset can generate molecular structures distinct from those seen during training, further validating its ability to foster innovation in molecular design.
Overall, the results above indicate that our dataset can significantly benefit molecule generation, offering a valuable resource for AI-driven molecule design and discovery.
4.3 Molecular Property Prediction Experiments
Accurate prediction of molecular properties is fundamental to drug discovery. To evaluate M3-20M’s effectiveness in property prediction tasks, we conduct extensive experiments following the settings in MoleculeNet wu2018moleculenet . Our evaluation encompasses regression and classification tasks, addressing diverse molecular properties crucial for drug development. We employ FP-ICL for regression tasks, both FP-ICL and parameter-efficient fine-tuning with LoRA for classification tasks. To prevent data leakage, we firstly partition datasets into non-overlapping training and test sets. Secondly, we mask all the target property-related information in textual descriptions during training and evaluation. We perform each experiment five times and report the averaged result and variance.
4.3.1 Regression Tasks
Experimental details. We apply FP-ICL for regression tasks on the QM9-MM dataset under three settings: using SMILES strings only, using both SMILES strings and 3D coordinates, and using SMILES strings with 3D coordinates plus textual descriptions. We focus on predicting dipole moment ($\mu$) and isotropic polarizability ($\alpha$). We use GPT-3.5, GPT-4, and GLM-4 APIs as LLMs, and randomly select four molecules from QM9-MM for the LLMs in the prompt according to the three specific settings.
Performance metrics. The metric we use for the regression tasks is Mean Absolute Error (MAE), measuring the average error between predicted values and actual values. The smaller the MAE value, the more accurate the prediction. Formally, MAE is evaluated as follows: